OpenStudy (karim728):

Logic: write the converse and contrapositive of the following statement: for all real numbers x, if -10 my answers: converse: for all real numbers x, if x+1>0 then -1x>=0

OpenStudy (karim728):

@triciaal @Nnesha

OpenStudy (karim728):

@zepdrix

zepdrix (zepdrix):

Converse is correct :) We have make some clever adjustments for the Contrapositive though.

zepdrix (zepdrix):

The problem we is that \(\large\rm x+1<0\) is not the negation of \(\large\rm x+1>0\)

OpenStudy (karim728):

x+1>=0 maybe

zepdrix (zepdrix):

You need to also include the 0 value (since you didn't at the start.). \(\large\rm \neg\left[x+1>0\right]\equiv \left[x+1\le0\right]\) Yes good good :)

zepdrix (zepdrix):

Oh maybe I misread your answer too quickly hehe

zepdrix (zepdrix):

We flip the sign, but we also include 0.

OpenStudy (karim728):

ooh so x+1>=0 would make it correct right

zepdrix (zepdrix):

No. You didn't flip your inequality sign. x+1 > 0 Should change to x+1 <= 0

zepdrix (zepdrix):

But we also need to fix the other inequality. That one is a bit more difficult.

zepdrix (zepdrix):

\[\large\rm -1<x\le0\]Drawing a number line will really help on this one.

OpenStudy (karim728):

ok then .. if i undertand correctly for all real numbers x, if x+1 <=0 then -1>=x>0

zepdrix (zepdrix):

Do you understand how I graphed the inequality? It's this shaded region in the middle for \(\large\rm -1<x\le0\) ya?